OpenStudy (anonymous):

the radius of the circle 2x² + 2y² - 3x + 4y - 1 = 0 a. √ 33 4 b. 33 16 c. √ 33 3 d. 17 please help me to get it. i need you solutions.

OpenStudy (phebe):

What do you think the answer is?

OpenStudy (anonymous):

b?

OpenStudy (loser66):

justify your answer. don't guess

OpenStudy (anonymous):

ok i don't know how to get the real answer. so please help me to know that.

OpenStudy (anonymous):

Remember the equation of a circle. (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

manipulate the equation so that it fits this format. Then you will see the radius

OpenStudy (loser66):

do you know how to make a perfect square?

OpenStudy (anonymous):

how can i apply the formula in the question?

OpenStudy (loser66):

ok, take steps, you do, I guide, deal?

OpenStudy (anonymous):

ok men.

OpenStudy (loser66):

if I re-arrange the given equation on this way, is it ok to you? \[2x^2-3x+2y^2+4y-1=0\]

OpenStudy (anonymous):

sure

OpenStudy (loser66):

now, I just take the x part only (don't worry, I will not let y free, just do later) \[2x^2 -3x\] to get this part become a perfect square, do you know how to do?

OpenStudy (anonymous):

no, i don't what will i do.

OpenStudy (anonymous):

but will try to listen just to understand

OpenStudy (anonymous):

x² - 3/2x = 0 x² - 3/4x = (- 3/4)² (x - 3/4)² = 9/16 x - 3/4 = 3/4 ?

OpenStudy (loser66):

nope, I am trying to make it as simple as possible to explain you. Wait, I will not let you alone

OpenStudy (anonymous):

ohh thank you so much bro! i'll appreciate your kindness.

OpenStudy (anonymous):

bro?

OpenStudy (loser66):

I am writing out, will scan and post after finish.

OpenStudy (anonymous):

ooh thanks men! I hope I can return the favor some day.

OpenStudy (loser66):

finally!!!

OpenStudy (loser66):

If you don't understand at somewhere, just make questions hihihi...

OpenStudy (anonymous):

so the final asnwer is 33 over 16?

OpenStudy (loser66):

\[\dfrac{\sqrt{33}}{4}\]

OpenStudy (anonymous):

thanks men!